Jay Leno’s collection of cars and motorcycles totals 187. When he checks the air in the tires he has 588 tires to check. How many cars and how many motorcycles does he own? Assume that the cars all have four tires and the motorcycles have two.

Answer by stanbon(75887)

Jay Leno’s collection of cars and motorcycles totals 187. When he checks the air in the tires he has 588 tires to check. How many cars and how many motorcycles does he own? Assume that the cars all have four tires and the motorcycles have two.
---------------
Vehicle Equation: c + m = 187
Tire Equation::::4c + 2m = 588
----------------------------------------
Multiply thru the vehicle equation by 2 to get:
2c + 2m = 2*187
---
Subtract that from the tire equation:
2c = 588-2*187
c = 107 (# of cars)
---------------------------
Substitute that into c + m = 187 to solve for "m":
107 + m = 187
m = 80 (# of motorcycles)
============================
Cheers,
Stan H.